OpenStudy (anonymous):

complete pattern 0, 1, 3, 7, 15, __,__

jimthompson5910 (jim_thompson5910):

hints: 1-0 = 1 3-1 = 2 7-3 = 4 15-7 = 8

OpenStudy (nirmalnema):

it is the series of geometric progression